### v1

# Dynamic Structural Causal Models

Abstract
We study a specific type of SCM, called a Dynamic Structural Causal Model (DSCM), whose endogenous variables represent functions of time, which is possibly cyclic and allows for latent confounding. As a motivating use-case, we show that certain systems of Stochastic Differential Equations (SDEs) can be appropriately represented with DSCMs. An immediate consequence of this construction is a graphical Markov property for systems of SDEs. We define a time-splitting operation, allowing us to analyse the concept of local independence (a notion of continuous-time Granger (non-)causality). We also define a subsampling operation, which returns a discrete-time DSCM, and which can be used for mathematical analysis of subsampled time-series. We give suggestions how DSCMs can be used for identification of the causal effect of time-dependent interventions, and how existing constraint-based causal discovery algorithms can be applied to time-series data.
1 Introduction
Many real-world systems exhibit a non-trivial dependence on time. In science and engineering, such systems are often modelled with Ordinary Differential Equations when the dynamics are deterministic, with Random Differential Equations (RDEs) when the dynamics of a population is described but every individual can have different (deterministic) dynamics, and with Stochastic Differential Equations (SDEs) when the dynamics of each individual are inherently stochastic (Banks et al., , 2014). In this work, we will focus on the latter formalism, which generalises the former two. SDEs do not come formally equipped with causal semantics. As their mathematical formulation is rather technical, it is a nontrivial task to apply existing methods of causal reasoning to this model class. In this work, we investigate how SDEs can be mapped into a Dynamic Structural Causal Models (DSCMs), originally introduced by Rubenstein et al., (2018) and refined in this work, to naturally bridge this gap. More precisely, we
  1. 1.
    refine the definition of DSCMs to make them specific versions of SCMs as defined by Bongers et al., (2021), enhancing them with a Markov property based on $\sigma$-separation;
  2. 2.
    show how certain uniquely solvable systems of SDEs can be transformed into DSCMs;
  3. 3.
    define time-splitting and subsampling as formal tools to analyse the time-dependent structure of DSCMs;
  4. 4.
    show how time-split DSCMs allow for a graphical interpretation of local independence and that the graph of ‘collapsed’ DSCMs are local independence graphs, yielding a local independence Markov property;
  5. 5.
    give further suggestions of how existing results for ‘static’ causal effect identification and constraint-based causal discovery are naturally applicable to DSCMs. [^1] All proofs are provided in Appendix C.
2 Systems of Stochastic Differential Equations
Consider the time-index $\mathcal{T}:=[0,T]$ for some $T\in\mathbb{R}$. Let $D(\mathcal{T},\mathbb{R}^{n})$ be the space of càdlàg222continue à droite, limité à gauche functions from $\mathcal{T}$ to $\mathbb{R}^{n}$. Throughout, we equip $D(\mathcal{T},\mathbb{R}^{n})$ with the $J_{1}$ topology (Skorokhod, , 1956) and the Borel sigma-algebra, making it a standard Borel space. Let $(\Omega,\mathbb{F},\mathbb{P})$ denote a filtered probability space with filtration $\mathbb{F}:=(\mathcal{F}_{t})_{t\in\mathcal{T}}$ that is right-continuous and contains all $\mathbb{P}$-null-sets. A process $X:\Omega\times\mathcal{T}\to\mathbb{R}^{n}$ is $\mathbb{F}$-adapted if $X(t)\in\mathcal{F}_{t}$ for all $t\in\mathcal{T}$; for predictable processes we have $X(t)\in\mathcal{F}_{t-}$, where $\mathcal{F}_{t-}:=\bigcap_{s<t}\mathcal{F}_{s}$. Let $\mathbb{S}_{\mathbb{F}}(\mathcal{T},\mathbb{R}^{n})$ denote the class of $\mathbb{R}^{n}$-valued semimartingales, i.e. processes $X:\Omega\times\mathcal{T}\to\mathbb{R}^{n}$ that are $\mathbb{F}$-adapted with càdlàg sample paths a.s., and with (unique) decomposition $X=A+M$ with $A$ a predictable process of finite variation and $M$ a local martingale. Let $\mathcal{P}_{b}$ denote the set of locally bounded predictably measurable processes. The stochastic integral of $G$ w.r.t. $H$
$J:\mathcal{P}_{b}\times\mathbb{S}\to\mathbb{S},~{}~{}(G,H)\mapsto\int_{0}^{t}G(s)\mathrm{d}H(s)$ (1)
is a well-defined notion (Protter, (2005), Thm. IV.15).
For a formal definition of a system of stochastic differential equations, we require the following notion:
Assumption 1.
Let $f:\mathcal{T}\times D(\mathcal{T},\mathbb{R}^{m})\to\mathbb{R}^{n}$ be a Borel-measurable function such that
  1. i)
    $t\mapsto f(t,x)\in D(\mathcal{T},\mathbb{R}^{n})$ for all $x\in D(\mathcal{T},\mathbb{R}^{m})$,
  2. ii)
    $f(t,x)=f(t,x^{t})$ for all $x\in D(\mathcal{T},\mathbb{R}^{m})$ and $t\in\mathcal{T}$, where $x^{t}(s):=x(s\wedge t)$.
Remark 1.
If $f$ satisfies Assumption 1 and $X\in\mathbb{S}_{\mathbb{F}}(\mathcal{T},\mathbb{R}^{m})$, then the process $f(t,X)$ is in $\mathbb{S}_{\mathbb{F}}(\mathcal{T},\mathbb{R}^{n})$, and $f(t-,X):=\lim_{s\uparrow t}f(t,X)$ is an element of $\mathcal{P}_{b}$ (Przybyłowicz et al., , 2023).
Definition 1 (System of SDEs, solutions).
Given a filtered probability space $(\Omega,\mathbb{F},\mathbb{P})$, a system of stochastic differential equations (SDEs) is a tuple $\mathcal{D}=\left\langle V,W,(X_{V}^{0},X_{W}),f\right\rangle$ with
  1. i)
    $V$ and $W$ finite disjoint index sets,
  2. ii)
    for each $v\in V$ a random variable $X_{v}^{0}:\Omega\to\mathbb{R}$, and for each $w\in W$ a stochastic process $X_{w}\in\mathbb{S}_{\mathbb{F}}(\mathcal{T},\mathbb{R})$,
  3. iii)
    for each $v\in V$ a stochastic differential equation
    $f_{v}:X_{v}(t)=X_{v}^{0}+\int_{0}^{t}g_{v}(s-,X_{\alpha(v)})\mathrm{d}h_{v}(s,X_{\beta(v)})$
    for all $t\in\mathcal{T}$, for sets $\alpha(v),\beta(v)\subseteq V\cup W$, and functions $g_{v}:\mathcal{T}\times D(\mathcal{T},\mathbb{R}^{|\alpha(v)|})\to\mathbb{R}^{m_{v}}$ and $h_{v}:\mathcal{T}\times D(\mathcal{T},\mathbb{R}^{|\beta(v)|})\to\mathbb{R}^{m_{v}}$ for some $m_{v}\in\mathbb{N}$, both satisfying Assumption 1.333We write $\int_{0}^{t}g\mathrm{d}h$ for $\sum_{i=1}^{m}\int_{0}^{t}g_{i}\mathrm{d}h_{i}$.
The processes $g_{v}(t,X_{\alpha(v)})$ and $h_{v}(t,X_{\beta(v)})$ are called integrands and integrators respectively. A solution [^4] Commonly referred to as a strong solution. of a system of SDEs $\mathcal{D}$ is a stochastic process $X_{V}\in\mathbb{S}_{\mathbb{F}}(\mathcal{T},\mathbb{R}^{|V|})$ such that $X_{V}=f_{V}(X_{V}^{0},X_{V},X_{W})$ holds $\mathbb{P}$-a.s.
By Remark 1, we have for semimartingales $X_{\alpha(v)}$ and $X_{\beta(v)}$ that $g_{v}(s-,X_{\alpha(v)})$ is in $\mathcal{P}_{b}$ and $h_{v}(s,X_{\beta(v)})$ is in $\mathbb{S}$, making the stochastic integrals well-defined (see equation 1). Note that we typically have $v\in\alpha(v)$. Definition 1 generalises common definitions of SDEs, which typically only model the relations between the variables that appear in the integrands, and assume the integrators to be given. We allow for explicit modelling of the variables that appear in the integrators because it is mathematically feasible, has applications in controlled SDEs (Lyons et al., , 2007), and allows for the modelling of instantaneous effects in time-split DSCMs (see Section 5).
Remark 2.
For specific choices of integrators and integrands, SDEs can model some special cases:
  1. 1.
    If $g(t,X_{\alpha})=g(t,X_{\alpha}(t))$ and $h$ is a Lévy process independent of $X^{0}$, then any solution $X$ is temporally Markov [^5] Process $X\in\mathbb{S}$ is called temporally Markov if $\mathbb{P}(X_{t+s}|\mathcal{F}_{t})=\mathbb{P}(X_{t+s}|X_{t})$ for all $s,t\in\mathcal{T}$ such that $s+t\in\mathcal{T}$; it is (temporally) strong Markov if this holds for any stopping time $t$. (Protter, , 2005, Theorem V.6.32). If additionally $g(t,X_{\alpha})=g(X_{\alpha}(t))$, then the solution is strong Markov; such dynamics are called time-invariant.
  2. 2.
    If $h(s)=s$, the ‘stochastic integral’ $\int g(s,X_{\alpha})\mathrm{d}h(s)$ reduces to the Riemann integral $\int g(s,X_{\alpha})\mathrm{d}s$, so a system of Random Differential Equations is a special case of a system of SDEs.
  3. 3.
    If $h(t,W)=W(t)$ is a Brownian Motion, any solution $X$ has continuous and possibly non-differentiable sample paths. If $h(t,W)=(t,W(t))$ and the dynamics are time-invariant, such an SDE is called an Itô diffusion.
  4. 4.
    If $h(t,N)=N(t)$ is a jump process (e.g. a Poisson process), the solution $X$ can have jumps as well. If $h(t,W,N)=(t,W(t),N(t))$ and the dynamics are time-invariant, such an SDE is called a jump-diffusion.
  5. 5.
    If $H$ is a process with differentiable sample paths, an RDE $X=\int g(H^{\prime}(s),X(s))\mathrm{d}s$ with $g$ linear in its first argument can be suitably modelled with an SDE $X=\int f(X)\mathrm{d}H(s)$ (Lyons et al., , 2007). The SDE formulation generalises this concept to a large class of non-differentiable processes $H$.
  6. 6.
    The functional relationship $X(t)=h(t,X_{\beta})$ can be equivalently expressed as $X(t)=h(0,X_{\beta})+\int_{0}^{t}\mathrm{d}h(s,X_{\beta})$.
Example 1.
Consider the following system of Stochastic Differential Equations
$\mathcal{D}:\begin{cases}X_{1}(t)=X_{1}^{0}+\int_{0}^{t}g_{1}(s-,X_{1},X_{3})\mathrm{d}W(s)\\ X_{2}(t)=X_{2}^{0}+\int_{0}^{t}g_{2}(s-,X_{1},X_{2})\mathrm{d}N(s)\\ X_{3}(t)=X_{3}^{0}+\int_{0}^{t}g_{3}(s-,X_{2},X_{3})\mathrm{d}W(s)\\ X_{4}(t)=X_{4}^{0}+\int_{0}^{t}g_{4}(s-,X_{4})\mathrm{d}X_{2}(s)\end{cases}$
with $W$ a Brownian motion, $N$ a Poisson process (independent from $W$), and $X_{1}^{0},X_{2}^{0},X_{3}^{0},X_{4}^{0}$ independent random variables. Note that $X_{2}$ is the integrator of $X_{4}$.
To ensure existence and uniqueness of solutions of a system of SDEs, we require additional regularity of the function $g$:
Assumption 2.
Let $g:\mathcal{T}\times D(\mathcal{T},\mathbb{R}^{m})\to\mathbb{R}^{n}$ satisfy Assumption 1, and let there exist a $K\in D(\mathcal{T},(0,\infty))$ such that for all $x,x_{1},x_{2}\in D(\mathcal{T},\mathbb{R}^{m})$ we have
$|g(t,x)|$ $\leq K(t)\left(1+\sup_{0\leq s\leq t}|x(s)|\right)$
$|g(t,x_{2})-g(t,x_{1})|$ $\leq K(t)\sup_{0\leq s\leq t}|x_{2}(s)-x_{1}(s)|.$
Theorem 1 (Przybyłowicz et al., (2023)).
For given functions $g:\mathcal{T}\times D(\mathcal{T},\mathbb{R}^{k+1})\to\mathbb{R}^{m}$ and $h:\mathcal{T}\times D(\mathcal{T},\mathbb{R}^{\ell})\to\mathbb{R}^{m}$ that satisfy Assumptions 2 and 1 respectively, there exists a Skorohod measurable mapping
$I:\mathbb{R}\times D(\mathcal{T},\mathbb{R}^{k})\times D(\mathcal{T},\mathbb{R}^{\ell})\to D(\mathcal{T},\mathbb{R})$
such that for any filtered probability space $(\Omega,\mathbb{F},\mathbb{P})$ with square-integrable random variable $X^{0}:\Omega\to\mathbb{R}$ and semimartingales $A\in\mathbb{S}_{\mathbb{F}}(\mathcal{T},\mathbb{R}^{k}),B\in\mathbb{S}_{\mathbb{F}}(\mathcal{T},\mathbb{R}^{\ell})$, the process $X:=I(X^{0},A,B)\in\mathbb{S}_{\mathbb{F}}(\mathcal{T},\mathbb{R})$ is the unique solution of
$X(t)=X^{0}+\int_{0}^{t}g(s-,X,A)\mathrm{d}h(s,B).$
We refer to such a solution function $I$ as an Itô map. The above result straightforwardly extends to processes $X$ taking values in $\mathbb{R}^{n}$.
2.1 Towards causal reasoning with SDEs
To analyse whether a system of SDEs can be solved, we use the dependency structure between the processes, which can be suitably depicted with a graph:
Definition 2 (Graph of system of SDEs).
For a system of SDEs $\mathcal{D}=\left\langle V,W,(X_{V}^{0},X_{W}),f\right\rangle$, let $G(\mathcal{D})=(V\cup W,E)$ be the directed graph where $E:=\{v\to w:w\in V,v\in\alpha(w)\cup\beta(w)\setminus\{w\}\}$.
Example 2.
Let $\mathcal{D}$ be the systems of SDEs from Example 1. The graph $G(\mathcal{D})$ is given in Figure 1. Outgoing edges of integrators are displayed in red.
[…]
Figure 1: The directed graph $G(\mathcal{D})$.
By partitioning the set of endogenous processes $V$ into strongly connected components (SCCs) (subsets whose every two vertices are connected by a directed path) and assuming that no element of such a SCC is the integrator of another element in the SCC (so no directed cycle contains a red edge), we can iteratively apply Theorem 1 along the topological order of the SCCs to obtain a unique solution.
Proposition 1 (Unique solvability of a system of SDEs).
[…]
We refer to a system of SDEs $\mathcal{D}$ that satisfies Proposition 1 as uniquely solvable.
Remark 3.
A system of SDEs is assumed to have a certain causal interpretation: for the SDEs in Definition 1, it is assumed that the LHS is determined by variables on the RHS. In Section 4 we will indeed interpret the Itô map $I_{v}$ for and SDE $f_{v}$ as a structural equation of a DSCM. This construction then gives a clear causal interpretation to the graph $G(\mathcal{D})$, where the processes $V$ and $W$ are interpreted as endogenous and exogenous respectively. Compared to existing work where for a systems of constraints one deduces the causal structure (e.g. Iwasaki and Simon, , 1994; Blom et al., , 2021), our setting makes relatively strong assumptions.
Another important causal aspect that is engrained in the system of SDEs is whether processes exhibit instantaneous dependencies. Although not explicitly stated in Przybyłowicz et al., (2023), one can verify via their proof of Theorem 1 that the Itô map $I$ satisfies $I(x^{0},a,b)^{t}=I(x^{0},a^{t-},b^{t})^{t}$. As it will play an important role in the analysis of temporal dynamics of DSCMs, we formally define this concept:
Definition 3 (Adapted, predictable dependence).
Let $g:D(\mathcal{T},\mathbb{R}^{m})\to D(\mathcal{T},\mathbb{R}^{n})$ be a Borel-measurable function. We say that the dependence of $g$ on $x$ is adapted if $g(x)^{t}=g(x^{t})^{t}$ and predictable if $g(x)^{t}=g(x^{t-})^{t}$ for all $t\in\mathcal{T}$.
Remark 4.
For given process $X$ and $Y:=g(X)$ for measurable $g$, we have $Y\in\mathcal{F}_{t}^{X}$ if $g$ is adapted, and $Y\in\mathcal{F}_{t-}^{X}$ if $g$ is predictable, [^7] Here, $\mathcal{F}_{t}^{X}$ denotes the filtration generated by $X$. motivating the nomenclature adapted and predictable (see Appendix A, Lemma 1). For any uniquely solvable system of SDEs $\mathcal{D}$ and any $v\in V$, the process $X_{v}$ has predictable dependence on $X_{\alpha(v)}$ and adapted (and thus possibly instantaneous) dependence on $X_{\beta(v)}$. Graphically, we depict adapted dependencies with red edges.
3 Dynamic Structural Causal Models
Following Bongers et al., (2021) and Forré and Mooij, (2023), a Structural Causal Model (SCM) is a tuple $\mathcal{M}=\left\langle V,W,\mathcal{X}_{V},\mathcal{X}_{W},f,\mathbb{P}(X_{W})\right\rangle$ where $V,W$ are disjoint finite index sets of endogenous variables and exogenous variables respectively, the domains $\mathcal{X}_{V}=\prod_{i\in V}\mathcal{X}_{i}$ and $\mathcal{X}_{W}=\prod_{i\in W}\mathcal{X}_{i}$ are products of standard Borel spaces $\mathcal{X}_{i}$, the exogenous distribution $\mathbb{P}(X_{W})=\bigotimes_{w\in W}\mathbb{P}(X_{w})$ is a product of probability distributions, and the causal mechanism $f:\mathcal{X}\to\mathcal{X}_{V}$ is a measurable function. Typically, SCMs are used to model ‘static’ systems, where variables take values in $\mathbb{R}$, for example. Discrete-time dynamical systems are often modelled with a copy of all variables $X_{V}$ for every time-step; an approach that does not directly apply to continuous-time dynamical systems. Inspired by (Rubenstein et al., , 2018) and Bongers et al., (2022), we define a specific type of SCM where variables are functions on some time-interval, and the structural equations respect temporal causality:
Definition 4 (DSCM).
Given a time-interval $\mathcal{T}=[0,T]$ for some $T\in\mathbb{R}_{+}$ or $\mathcal{T}\subseteq\{1,...,T\}$ for some $T\in\mathbb{N}$, a Dynamic Structural Causal Model is an SCM $\mathcal{M}=\left\langle V,W,\mathcal{X}_{V},\mathcal{X}_{W},f,\mathbb{P}\right\rangle$ with
  1. i)
    $V:=V_{p}\times\mathcal{E}$ with endogenous processes $V_{p}$ and $\mathcal{E}$ a set of disjoint evaluation intervals or points in $\mathcal{T}$; [^8] For $(v,\mathcal{I})\in V\times\mathcal{E}$ we write $v^{\mathcal{I}}:=(v,\mathcal{I})$ and $X_{v}^{\mathcal{I}}:=X_{(v,\mathcal{I})}$ and $X_{v}:=X_{v}^{\mathcal{T}}$.
  2. ii)
    exogenous processes $W$;
  3. iii)
    standard Borel spaces $\mathcal{X}_{V}=\bigtimes_{(v,\mathcal{I})\in V\times\mathcal{E}}D(\mathcal{I},\mathbb{R})$ and $\mathcal{X}_{W}=\bigtimes_{w\in W}D(\mathcal{T},\mathbb{R})$;
  4. iv)
    measurable structural equations $f:\mathcal{X}_{V}\times\mathcal{X}_{W}\to\mathcal{X}_{V}$ that are adapted (see Definition 3);
  5. v)
    exogenous distribution $\mathbb{P}(X_{W})=\bigotimes_{w\in W}\mathbb{P}(X_{w})$.
The set $\mathcal{E}$ denotes the time intervals (or points) on which we evaluate the processes $X_{V_{p}}$. For example, when ={{0},($\mathcal{E}=\{\{0\},(0,T]\}$, for every $v\in V_{p}$ we will consider the variables $X_{v}^{0}$ and $X_{v}^{(0,T]}$. We refer to a DSCM as collapsed if $\mathcal{E}=\mathcal{T}$. For $v\in V_{p}$, we denote with $\alpha(v)$ and $\beta(v)$ the sets of parents of $v$ on which it has predictable and adapted dependence respectively.
For notational simplicity, endogenous parameters like initial values or parameters that ‘materialise’ at some $t_{0}\in\mathcal{T}$ (i.e. other variables are only allowed to depend on it from $t_{0}$ onwards) are modelled as elements of $D(\{t_{0}\},\mathbb{R})$. For a function $x:(s,t)\to\mathbb{R}$ we use the convention that $x(r)=x(s+)$ for all $r<s$ and $x(u)=x(t-)$ for all $u>t$ [^9] Here, $x(t+)$ denotes $\lim_{s\downarrow t}x(s)$.
We have refined the definition of DSCMs as given by Rubenstein et al., (2018) by 1) ensuring the structural equations respect temporal causality, 2) allowing for stochasticity through the exogenous variables, and 3) picking a standard Borel space ($D(\mathcal{T},\mathbb{R})$) as space of trajectories, making for all $A,B\subseteq V\cup W$ the conditional distribution $\mathbb{P}(X_{A}\mid X_{B})$ well-defined, and with that giving a well-defined notion of conditional independence. DSCMs differ from Structural Dynamical Causal Models (SDCMs) (Bongers et al., , 2022) as we model entire sample paths as a whole, instead of separately at every time $t\in\mathcal{T}$. When $\mathcal{T}$ or $\bigcup\mathcal{E}$ is a discrete set, the DSCM reduces to a structural vector autoregression model (Peters et al., , 2013; Malinsky and Spirtes, , 2018) with possible serial dependence of the noise variables, or a Dynamic Bayesian Network (Dean and Kanazawa, , 1989).
Perfect interventions on DSCMs are defined exactly the same as for SCMs:
Definition 5 (Perfect interventions on DSCMs).
Given a DSCM $\mathcal{M}$, intervention target $T\subseteq V$ and intervention value $x_{T}\in\mathcal{X}_{T}$, the perfectly intervened DSCM is defined as $\mathcal{M}_{\operatorname*{\mathrm{do}}(X_{T}=x_{T})}:=\left\langle V,W,\mathcal{X}_{V},\mathcal{X}_{W},f^{\circ},\mathbb{P}(X_{W})\right\rangle$ where the components of the intervened causal mechanism $f^{\circ}:\mathcal{X}_{V}\times\mathcal{X}_{W}\to\mathcal{X}_{V}$ are given by
$f^{\circ}_{v}(x_{V},x_{W}):=\begin{cases}f_{v}(x_{V},x_{W})&\text{if $v\in V\setminus T$}\\ x_{v}&\text{if $v\in T$}.\end{cases}$
For a given SCM $\mathcal{M}$, a random variable $X_{V}\in\mathcal{X}_{V}$ is a solution of the SCM if $X_{V}=f(X_{V},X_{W})$ holds $\mathbb{P}(X_{W})$-a.s. Existence and uniqueness of solutions are generally not guaranteed for (intervened) cyclic SCMs. To this end, Bongers et al., (2022) introduced the following class of SCMs:
Definition 6 (Simple (D)SCM).
A (D)SCM is simple if for all $O\subseteq V$ (write $T:=V\setminus O$) there exists a unique (up to modifications [^10] For every $x_{T}$, $g_{O}(x_{T},x_{W})$ is $\mathbb{P}(X_{W})$ a.e. unique.) measurable function $g_{O}:\mathcal{X}_{T}\times\mathcal{X}_{W}\to\mathcal{X}_{O}$ such that for all $x_{T}\in\mathcal{X}_{T}$ and $\mathbb{P}(X_{W})$-almost all $x_{W}\in\mathcal{X}_{W}$ we have $g_{O}(x_{T},x_{W})=f_{O}(x_{T},g_{O}(x_{T},x_{W}),x_{W})$.
For a simple DSCM $\mathcal{M}$, the random variable $g_{O}(x_{T},X_{W})$ is a solution for $\mathcal{M}_{\operatorname*{\mathrm{do}}(X_{T}=x_{T})}$, implying the existence of a well-defined interventional distribution $\mathbb{P}(X_{V}|\operatorname*{\mathrm{do}}(X_{T}=x_{T}))$. The function $g_{O}$ is referred to as a solution function for $\mathcal{M}_{\operatorname*{\mathrm{do}}(X_{T}=x_{T})}$. The following proposition shows that solutions of DSCMs obey temporal causality:
Proposition 2.
Given a simple DSCM $\mathcal{M}$, the solution function $g_{O}(x_{T},x_{W})$ of the intervened DSCM $\mathcal{M}_{\operatorname*{\mathrm{do}}(X_{T}=x_{T})}$ is adapted.
Remark 5.
Following Remark 4, adaptedness of the solution function $g_{O}(x_{T},x_{W})$ is equivalent to adaptedness of solution process $X_{V}:=g_{O}(x_{T},X_{W})$ to a canonical filtration on the underlying probability space $\mathcal{X}_{T}\times\mathcal{X}_{W}$. For more information, see Appendix A.
In line with marginalisation of probability distributions, the marginalisation of a simple DSCM is defined as follows:
Definition 7 (Marginalisation).
For simple DSCM $\mathcal{M}$ and subset $L\subseteq V$ (write $O:=V\setminus L$), the marginalised DSCM $\mathcal{M}_{\operatorname*{\mathrm{marg}}(L)}:=\left\langle O,W,\mathcal{X}_{O},\mathcal{X}_{W},\tilde{f},\mathbb{P}(X_{W})\right\rangle$ has marginal causal mechanism
$\tilde{f}(x_{O},x_{W}):=f_{O}(x_{O},g_{L}(x_{O},x_{W}),x_{W}),$
where $g_{L}$ is the solution function in $\mathcal{M}_{\operatorname*{\mathrm{do}}(X_{O}=x_{O})}$.
Proposition 3.
The class of simple DSCMs is closed under marginalisation.
Bongers et al., (2021) show that marginalised simple SCMs are observationally, interventionally and counterfactually equivalent to the original SCM with respect to the observed variables. This makes marginalisation a powerful notion of abstraction, as any causal inference on the observed part of a system remains valid in the underlying system.
3.1 A Markov property for simple DSCMs
For SCM $\mathcal{M}$ and nodes $i\in V\cup W$ and $j\in V$ we call $i$ a parent of $j$ if there does not exist a measurable function $\tilde{f}_{j}:\mathcal{X}_{V\cup W\setminus\{i\}}\to\mathcal{X}_{j}$ such that $f_{j}(x_{V},x_{W})=\tilde{f}_{j}(x_{V\cup W\setminus\{i\}})$ for $\mathbb{P}(X_{W})$-almost all $x_{W}\in\mathcal{X}_{W}$ and for all $x_{V}\in\mathcal{X}_{V}$. We call a directed graph $G^{+}(\mathcal{M}):=(V\cup W,E)$ with directed edges $E=\{i\to j:i\in V\cup W,j\in V,\text{$i$ is a parent of $j$ in $\mathcal{M}$}\}$ the augmented graph of $\mathcal{M}$. Often we don’t display the exogenous nodes in the graph, in which case we consider the directed mixed graph $G(\mathcal{M})=(V,E^{\prime})$ where $E^{\prime}=E\cup\{i\leftrightarrow j:i,j\in V,(i\leftarrow k\to j)\in G(\mathcal{M})\text{ for some $k\in W$}\}$, where $E$ stems from $G^{+}(\mathcal{M})$. We call an SCM cyclic if its graph $G(\mathcal{M})$ has a directed cycle. Simple SCMs can be cyclic, but cannot have self-cycles. A suitable extension of $d$-separation (Pearl, , 2009) from DAGs to DMGs is the following notion of $\sigma$-separation:
Definition 8 ($\sigma$-separation, Forré and Mooij, (2017); Bongers et al., (2021)).
Given a DMG $G=(V,E)$, a set of nodes $C\subset V$ and a walk $\pi$ in DMG $G$:
  1. 1.
    a non-collider is called blockable if it has a child on the walk that is not in the same strongly connected component,
  2. 2.
    the walk $\pi$ is called $\sigma$-blocked by $C$ if there is a collider on $\pi$ that is not in $\operatorname*{\mathrm{Anc}}(C)$, or if there is a blockable non-collider on $\pi$ in $C$.
For sets of nodes $A,B,C\subseteq V$, we call $A$ and $B$ $\sigma$-separated given $C$, written $A\perp_{G}^{\sigma}B\operatorname*{|}C$, if all paths between $A$ and $B$ are $\sigma$-blocked by $C$.
In general $\sigma$-separation implies $d$-separation, and for acyclic DMGs the two notions coincide.
Example 3.
In the DMG $G$ from Figure 1 we have $X_{1}^{0}\operatorname*{\perp}^{d}_{G}X_{2}^{0}\mid X_{1},X_{2}$, and $X_{1}^{0}\operatorname*{\not\perp}^{\sigma}_{G}X_{2}^{0}\mid X_{1},X_{2}$.
Theorem 2 (Markov Property, Forré and Mooij, (2017); Bongers et al., (2021)).
For simple (D)SCM $\mathcal{M}$ with distribution $\mathbb{P}(X_{V},X_{W})$, graph $G(\mathcal{M})$ and (not necessarily disjoint) sets $A,B,C\subseteq V\cup W$, we have
$A\perp_{G(\mathcal{M})}^{\sigma}B\operatorname*{|}C\implies X_{A}\operatorname*{\perp\mkern-11.0mu\perp}_{\mathbb{P}}X_{B}\operatorname*{|}X_{C}.$
A simple (D)SCM $\mathcal{M}$ is called $\sigma$-faithful if the reverse implication holds; this need not hold in general.
4 Constructing a DSCM from a system of SDEs
As mentioned before, one of our goals is to map a system of SDEs $\mathcal{D}$ to a (D)SCM, so we can apply all tools that are available for SCMs to SDEs. However, an SDE cannot directly be interpreted as structural equation of a DSCM, as SDEs are not equations between variables taking values in $D(\mathcal{T},\mathbb{R})$, but in $\mathbb{S}(\mathcal{T},\mathbb{R})$. However, when $\mathcal{D}$ is uniquely solvable, we can unambiguously interpret the Itô maps of $\mathcal{D}$ as the structural equations of a DSCM $\mathcal{M}_{\mathcal{D}}$.
Definition 9 (DSCM induced by a system of SDEs).
Given a uniquely solvable system of SDEs $\mathcal{D}=\left\langle V_{p},W_{p},(X_{V_{p}}^{0},X_{W_{p}}),f\right\rangle$ on a filtered probability space $(\Omega,\mathbb{F},\mathbb{P})$ and time-interval $[0,T]$, we define the simple DSCM induced by $\mathcal{D}$ as $\mathcal{M}_{\mathcal{D}}:=\left\langle V,W,\mathcal{X}_{V},\mathcal{X}_{W},f^{*},\mathbb{P}(X_{W})\right\rangle$ with
  1. 1.
    $V:=V_{p}\times\mathcal{E}$ with $\mathcal{E}:=\mathcal{T}$;
  2. 2.
    $W:=W_{0}\cup W_{p}$ with $W_{0}:=\{v^{0}:v\in V_{p}\}$ a copy of $V_{p}$ as the indices of the initial values;
  3. 3.
    for every $v\in V_{p}$ the structural equation
    $f^{*}_{v}(x_{V},x_{W}):=I_{v}(x_{v}^{0},x_{\alpha(v)},x_{\beta(v)})$
    where $I_{v}$ is the Itô map for the SDE $f_{v}$;
  4. 4.
    $\mathcal{X}_{V}:=D(\mathcal{T},\mathbb{R}^{|V|})$ and $\mathcal{X}_{W}:=D(\mathcal{T},\mathbb{R}^{|W|})$;
  5. 5.
    $\mathbb{P}(X_{W})$ is the law of $(X_{V_{p}}^{0},X_{W_{p}})$.
Note that $G^{+}(\mathcal{M}_{\mathcal{D}})=G(\mathcal{D})$, and does not have any self-loops $v\to v$. Interventions on a SDE $\mathcal{D}$ can be defined similarly as for DSCMs, intervening and mapping to a DSCM commute, and $\mathcal{M}_{\mathcal{D}}$ is indeed simple (see Appendix B).
Example 4.
Let $\mathcal{D}$ be the systems of SDEs from Example 1. The structural equations of the induced DSCM $\mathcal{M}_{\mathcal{D}}$ are given by:
$\mathcal{M}_{\mathcal{D}}:\begin{cases}X_{1}=f_{1}(X^{0}_{1},X_{3},W)\\ X_{2}=f_{2}(X^{0}_{2},X_{1},N)\\ X_{3}=f_{3}(X^{0}_{3},X_{2},W)\\ X_{4}=f_{4}(X^{0}_{4},X_{2}).\end{cases}$
The augmented graph $G^{+}(\mathcal{M}_{\mathcal{D}})=G(\mathcal{D})$ is given in Figure 1, and $G(\mathcal{M}_{\mathcal{D}})$ is given in Figure 2, where the bidirected edge represents the latent confounder $W$.
[…]
Figure 2: The directed mixed graph $G(\mathcal{M}_{\mathcal{D}})$.
Remark 6.
A technical but important detail is that the Itô maps $f^{*}_{v}$ (and thus the structural equations of $\mathcal{M}_{\mathcal{D}}$) are well-defined everywhere on their domain, due to the result of Przybyłowicz et al., (2023). If one obtains such an Itô map through the factorisation lemma (Kallenberg, (2021), Lemma 1.14) it is not everywhere uniquely defined, rendering the effects of interventions outside this essential support ambiguous. It is also important to note that the Itô maps have the right measurability properties (with respect to the Borel sets generated by the $J_{1}$ topology).
Remark 7.
The Markov property for DSCMs (Theorem 2) is closely related to the Markov property for the SDCM framework for RDEs (Bongers et al., , 2022). We conjecture that for a given system $\mathcal{D}$ of uniquely solvable RDEs, the graph $G(\mathcal{M}_{\mathcal{D}})$ of DSCM $\mathcal{M}_{\mathcal{D}}$ is a subgraph of the graph $G(\mathcal{R}_{\mathcal{D}})$ of the SDCM $\mathcal{R}_{\mathcal{D}}$, which would make the DSCM Markov property at least as powerful as the SDCM Markov property.
5 Time-evaluations of DSCMs
[…]
Figure 3: The graph $G(\mathcal{M})$, time-split $G(\mathcal{M}_{\operatorname*{ev}^{+}(0,s,t)})$ and subsampled $G(\mathcal{M}_{\operatorname*{ev}(0,s,t)})$.
Given a system of SDEs $\mathcal{D}$, we can construct a system of SDEs $\mathcal{D}_{\operatorname*{ev}(r)}$ by splitting up the time interval $[0,T]=[0,r)\cup\{r\}\cup(r,T]$, and consider the SDEs
$X_{v}(t)$ $=X_{v}^{0}+\int_{0}^{t}g_{v}(s-,X_{\alpha(v)})\mathrm{d}h_{v}(s,X_{\beta(v)})$
$X_{v}(r)$ $=X_{v}^{0}+\int_{0}^{r}g_{v}(s-,X_{\alpha(v)})\mathrm{d}h_{v}(s,X_{\beta(v)})$
$X_{v}(u)$ $=X_{v}(r)+\int_{r}^{u}g_{v}(s-,X_{\alpha(v)})\mathrm{d}h_{v}(s,X_{\beta(v)})$
for all $v\in V$ and all $t\in[0,r)$ and $u\in(r,T]$. We can do this type of time-splitting also on the level of a DSCM, to be formally defined below. The graphs of an example DSCM $\mathcal{M}$ and its time-split version $\mathcal{M}_{\operatorname*{ev}^{+}(0,s,t)}$ and subsampled version $\mathcal{M}_{\operatorname*{ev}(0,s,t)}$ are given in Figure 3.
Definition 10 (time-split DSCM).
Let a set of distinct time-indices $\tau=\{t_{1},...,t_{m}\}\subseteq\mathcal{T}$ and DSCM $\mathcal{M}=\left\langle V_{p}\times\mathcal{E},W,\mathcal{X}_{V},\mathcal{X}_{W},f,\mathbb{P}(X_{W})\right\rangle$ be given. The time-split DSCM $\mathcal{M}_{\operatorname*{ev}^{+}(\tau)}$ is defined as $\mathcal{M}_{\operatorname*{ev}^{+}(\tau)}=\left\langle\tilde{V},W,\mathcal{X}_{\tilde{V}},\mathcal{X}_{W},\tilde{f},\mathbb{P}(X_{W})\right\rangle$ with as differences from $\mathcal{M}$:
  1. i)
    $\tilde{V}=V_{p}\times\tilde{\mathcal{E}}$, where $\tilde{\mathcal{E}}$ is a further subdivision of $\mathcal{E}$, where for every $t\in\tau$, an interval $[s,u]\in\mathcal{E}$ containing $t$ is partitioned into $\{[s,t),\{t\},(t,u]\}\subseteq\tilde{\mathcal{E}}$; [^11] A similar construction applies for (half-)open intervals.
  2. ii)
    standard Borel space $\mathcal{X}_{\tilde{V}}=\bigtimes_{(v,\mathcal{I})\in\tilde{V}}D(\mathcal{I},\mathbb{R})$;
  3. iii)
    for every $(v,\mathcal{I})\in\tilde{V}$ a structural equation $\tilde{f}_{v^{\mathcal{I}}}:\mathcal{X}_{\tilde{V}}\times\mathcal{X}_{W}\to\mathcal{X}_{v^{\mathcal{I}}}$ with for all $t\in\mathcal{I}$
    $\tilde{f}_{v^{\mathcal{I}}}(x_{\tilde{V}},x_{W})(t)$ $:=f_{v}\left(x_{\alpha(v)}^{\tilde{\mathcal{E}}_{<\mathcal{I}}},x_{\beta(v)}^{\tilde{\mathcal{E}}_{\leq\mathcal{I}}},x_{W}\right)(t),$
    where ~:={~$\tilde{\mathcal{E}}_{\leq\mathcal{I}}:=\{\mathcal{I}^{\prime}\in\tilde{\mathcal{E}}\mid\exists s\in\mathcal{I}^{\prime},t\in\mathcal{I}:s\leq t\}$.
The above definition can straightforwardly be extended to differing time-evaluations $\tau_{v}$ for each $v\in V$. Note the explicit modelling of the predictable dependence on $\alpha$, and adapted dependence on $\beta$.
Definition 11 (Subsampled DSCM).
Given a DSCM $\mathcal{M}$, consider the time-split DSCM $\mathcal{M}_{\operatorname*{ev}^{+}(\tau)}$ with endogenous variables $V\times\mathcal{E}$. Let $\mathcal{E}^{\prime}\subseteq\mathcal{E}$ denote the set of non-singleton intervals. The subsampled DSCM is defined as $\mathcal{M}_{\operatorname*{ev}(\tau)}:=(\mathcal{M}_{\operatorname*{ev}^{+}(\tau)})_{\operatorname*{\mathrm{marg}}(V\times\mathcal{E}^{\prime})}$.
Example 5.
Let $\mathcal{M}_{\mathcal{D}}$ be the DSCM from Example 4, and let $\mathcal{M}:=(\mathcal{M}_{\mathcal{D}})_{\operatorname*{\mathrm{marg}}(L)}$ be the marginalisation of $\mathcal{M}_{\mathcal{D}}$ with respect to $L:=\{X_{3}^{0},X_{3},X_{4}^{0}\}$. The graphs $G(\mathcal{M})$, time-split $G(\mathcal{M}_{\operatorname*{ev}^{+}(s,t)})$ and subsampled $G(\mathcal{M}_{\operatorname*{ev}(s,t)})$ are depicted in Figure 3, where we assume all processes to be temporally Markov. We assume the processes $X_{1},X_{2}$ and $X_{4}$ to be temporally Markov for graphical appeal.
Definition 12 (Path concatenation).
Let $\mathcal{I}_{X},\mathcal{I}_{Y}$ be two adjacent intervals in $[0,T]$ with $s:=\sup\mathcal{I}_{X}=\inf\mathcal{I}_{Y}$, and let $X\in D(\mathcal{I}_{X},\mathbb{R})$ and $Y\in D(\mathcal{I}_{Y},\mathbb{R})$. The concatenation of $X$ and $Y$ is the path $X*Y\in D(\mathcal{I}_{X}\cup\mathcal{I}_{Y},\mathbb{R})$ defined by
$(X*Y)(t):=\begin{cases}X(t)&\text{if $t\in[0,s)\cap\mathcal{I}_{X}$}\\ Y(t+)&\text{if $t\in[s,T]\cap\mathcal{I}_{Y}$.}\end{cases}$
Definition 13 (Collapsed DSCM).
For a DSCM $\mathcal{M}=\left\langle V\times\mathcal{E},W,\mathcal{X}_{V},\mathcal{X}_{W},f,\mathbb{P}(X_{W})\right\rangle$, the collapsed DSCM is defined as $\mathcal{M}_{*}:=\left\langle\tilde{V},W,\mathcal{X}_{\tilde{V}},\mathcal{X}_{W},\tilde{f},\mathbb{P}(X_{W})\right\rangle$, where $\tilde{V}:=V\times\tilde{\mathcal{E}}$ with $\tilde{\mathcal{E}}:=\bigcup\mathcal{E}$ and structural equations
[…]
Note that this is consistent with the nomenclature that is earlier introduced, where we referred to a DSCM as collapsed if $\mathcal{E}=\mathcal{T}$.
The following proposition shows that no information is lost when time-splitting a DSCM, but not when subsampling a DSCM.
Proposition 4.
For a DSCM $\mathcal{M}$ with evaluation index $\mathcal{E}$ that partitions $\mathcal{T}$ and a sequence of increasing time-indices $\tau=(t_{1},...,t_{m})\subseteq\mathcal{T}$, we have $\mathcal{M}=\mathcal{M}_{\operatorname*{ev}^{+}(\tau)*}$. If $\tau\neq\mathcal{T}$, we have $\mathcal{M}\neq\mathcal{M}_{\operatorname*{ev}(\tau)*}$.
5.1 Local Conditional Independence
Throughout this section, let $\mathcal{M}$ be a simple DSCM with time-interval $\mathcal{T}=[0,T]$ for some $T\in\mathbb{R}$, with endogenous variables $A,B,C\in V$ such that $X_{A},X_{B},X_{C}\in\mathbb{S}_{\mathbb{F}_{\mathcal{X}_{W}}}$ (i.e. they are semimartingales). With respect to $\mathcal{F}^{A,B,C}$, such processes have a unique Doob-Meyer decomposition $X_{B}=\Lambda_{B}+M_{B}$, where ‘the drift term’ $\Lambda_{B}$ is a predictable $\mathcal{F}^{A,B,C}_{t}$-adapted process, and ‘the noise term’ $M_{B}$ is a $\mathcal{F}^{A,B,C}_{t}$-adapted martingale. We assume that the process $\Lambda_{B}$ is absolutely continuous, i.e. $\Lambda_{B}=\int_{0}^{t}\lambda_{B}(s)\mathrm{d}s$ for some $\mathcal{F}^{A,B,C}_{t}$-predictable intensity process $\lambda_{B}$.
Definition 14 (Local conditional independence).
We say that $X_{B}$ is locally conditionally independent from $X_{A}$ given $X_{C}$, written $X_{A}\not\to X_{B}\mid X_{C}$, if the intensity process $\lambda_{B}$ has a $\mathcal{F}_{t}^{B,C}$-predictable version. [^12] If one does not want to assume the existence of the intensity $\lambda_{B}$, local conditional independence can be defined more generally as $\Lambda_{B}$ having a $\mathcal{F}^{B,C}_{t}$-adapted version, or equivalently, that $M_{B}$ has a version that is a $\mathcal{F}^{B,C}_{t}$-adapted martingale (Mogensen and Hansen, , 2020).
In the following theorem, we relate local independence to statistical relation between the variables in time-split DSCMs.
Theorem 3.
Considering DSCM $\mathcal{M}$ and for every $t\in\mathcal{T}$ the time-split DSCM $\mathcal{M}_{\operatorname*{ev}^{+}(t)}$, we have that
$X_{B}^{t}\operatorname*{\perp\mkern-11.0mu\perp}X_{A}^{[0,t)}|X_{B}^{[0,t)},X_{C}^{[0,t)}\text{ for all $t\in\mathcal{T}$}$
$\implies$ $X_{A}\not\to X_{B}\mid X_{C}$
$\implies$ $\mathbb{E}[X_{B}^{t}\mid X_{A}^{[0,t)},X_{B}^{[0,t)},X_{C}^{[0,t)}]=\mathbb{E}[X_{B}^{t}\mid X_{B}^{[0,t)},X_{C}^{[0,t)}]$
for all $t\in\mathcal{T}$.
Definition 15 (Local independence graph, Mogensen and Hansen, (2022)).
Given a set $V$ of semimartingales $X_{v}$ with joint distribution $\mathbb{P}(X_{V})$, a directed mixed graph $G=(V,E)$ is a local independence graph for $\mathbb{P}(X_{V})$ if $A\to B$ not in $G(\mathcal{M})$ implies $X_{A}\not\to_{\mathbb{P}}X_{B}|X_{V\setminus A}$ for all $A,B\in V$.
Existing work on local independence often assume some kind of autonomy of the processes $X_{A},X_{B},X_{C}$. Aalen et al., (2012) assumes that the martingales $M_{A},M_{B},M_{C}$ are strongly orthogonal (i.e. so have zero quadratic covariation). Didelez, (2008); Røysland et al., (2024) consider pure jump-type processes and assume orthogonality, ensuring they never jump simultaneously. Christgau et al., (2023) require $X_{A},X_{C}$ to be predictable, which ensures $M_{A}=M_{C}=0$. We consider a slightly stronger notion of autonomy:
Assumption 3 (Independent integrators).
Let $\mathcal{M}$ be a collapsed DSCM with $\beta(v)\subseteq W$ and $\beta(v)\cap\beta(v^{\prime})=\emptyset$ for all $v,v^{\prime}\in V$.
In words, this means that there can be no instantaneous effects between the endogenous variables, nor any instantaneous dependencies between them through latent confounders.
Theorem 4.
For a collapsed DSCM $\mathcal{M}$ that satisfies Assumption 3 we have:
  1. 1.
    $G(\mathcal{M})$ is a local independence graph,
  2. 2.
    a $\sigma$-separation local independence global Markov property:
    $A\operatorname*{\perp}_{G(\mathcal{M})}^{\sigma}B|C\implies X_{A}\not\to X_{B}\mid X_{C}.$
Example 6.
The graph $G(\mathcal{M}_{\mathcal{D}})$ in Figure 2 is not guaranteed to be a local independence graph, but $G((\mathcal{M}_{\mathcal{D}})_{\operatorname*{\mathrm{marg}}(L)})$ with $L:=\{X_{3},X_{4}\}$ is.
We note that the above Markov property is weaker than existing Markov properties for $\delta$ or $\mu$ separation (e.g. Didelez, (2008) for orthogonal counting processes, and Mogensen et al., (2018) for Itô processes with independent noise), as $\delta$- and $\mu$-separation imply $\sigma$-separation. However, it applies to any simple DSCM induced by uniquely solvable systems of SDEs with independent integrators, so spans a much larger class of processes. [^13] A notable exception is given in Mogensen and Hansen, (2022) for Ornstein-Uhlenbeck processes with correlated noise.
6 Further Applications
In this section, we briefly elaborate how existing results for causal effect identification and causal discovery can directly be applied to DSCMs.
6.1 Causal Effect Identification
We note that perfect interventions can be modelled for an entire trajectory $X_{v}^{[0,T]}$ in a collapsed DSCM, on a specific subinterval $X_{v}^{[s,t]}$ with $[s,t]\subseteq\mathcal{T}$ in a time-split DSCM, or on a time-point $X_{v}^{t}$ in either a time-split or subsampled DSCM, e.g. in the DSCMs depicted in Figure 3.
The rules of do-calculus are valid for simple (D)SCMs (Forré and Mooij, , 2020). These criteria are stated in terms of so-called intervention nodes, which for variable $X\in V$ are defined as vertices $I_{X}$ in $G(\mathcal{M})$ with $I_{X}\to X$ as their only connected edge.
Theorem 5 (Do-calculus, Forré and Mooij, (2020)).
Given a (D)SCM $\mathcal{M}$ with $X,Y,Z,W\subseteq V$, we have
  • Insertion/deletion of observation:
    $Y\operatorname*{\perp}^{\sigma}X\mid Z,\operatorname*{\mathrm{do}}(W)\implies$
         $\mathbb{P}(Y\mid X,Z,\operatorname*{\mathrm{do}}(W))=\mathbb{P}(Y\mid Z,\operatorname*{\mathrm{do}}(W))$
  • Action/observation exchange:
    $Y\operatorname*{\perp}^{\sigma}I_{X}\mid X,Z,\operatorname*{\mathrm{do}}(W)\implies$
         $\mathbb{P}(Y\mid\operatorname*{\mathrm{do}}(X),Z,\operatorname*{\mathrm{do}}(W))=\mathbb{P}(Y\mid X,Z,\operatorname*{\mathrm{do}}(W))$
  • Insertion/deletion of actions:
    $Y\operatorname*{\perp}^{\sigma}I_{X}\mid Z,\operatorname*{\mathrm{do}}(W)\implies$
         $\mathbb{P}(Y\mid\operatorname*{\mathrm{do}}(X),Z,\operatorname*{\mathrm{do}}(W))=\mathbb{P}(Y\mid Z,\operatorname*{\mathrm{do}}(W))$.
As a consequence, generalised adjustment formulae (like backdoor adjustment) based on $\sigma$-separation criteria are also valid for simple DSCMs, as is the ID algorithm. For more information, we refer to Forré and Mooij, (2020).
The above mentioned results consider the identification of certain estimands in terms of observational distributions. We note that the construction of estimators for such expressions can be highly nontrivial when the variables take values in function spaces.
6.2 Causal Discovery
For a given simple DSCM with graph $G$, let $\textrm{IM}_{\sigma}(G)$ denote the $\sigma$-independence model of $G$, i.e. the set of all $\sigma$-separations that are present in $G$. Constraint-based causal discovery algorithms are methods that reconstruct (an equivalence class of) $G$ from the independence model $\textrm{IM}_{\sigma}(G)$. An example is Fast Causal Discovery (FCI) (Spirtes et al., , 1995; Zhang, , 2008), which outputs a Partial Ancestral Graph (PAG), and which is shown to be sound and complete for simple (and thus possibly cyclic) SCMs. In the following let $F$ denote FCI as a mapping from an independence model to a PAG.
Theorem 6 (FCI, Mooij and Claassen, (2020)).
Let $\mathcal{M}$ be a simple (D)SCM with graph $G$, then
  1. 1.
    FCI is sound: $G\in F(\textrm{IM}_{\sigma}(G))$
  2. 2.
    FCI is complete: let $G^{\prime}$ be another DMG, then $F(\textrm{IM}_{\sigma}(G))=F(\textrm{IM}_{\sigma}(G^{\prime}))$ iff $\textrm{IM}_{\sigma}(G)=\textrm{IM}_{\sigma}(G^{\prime})$.
Other constraint-based causal discovery algorithms that are known to be sound for (possibly cyclic) simple DSCMs are Local Causal Discovery (Cooper, , 1997; Mooij et al., , 2020) and Y-structures (Mani, , 2006; Mooij et al., , 2020).
In practice, one requires a conditional independence (CI) test to map a dataset with samples of an SCM $\mathcal{M}$ to an independence model (assuming faithfulness). If variables $X,Y,Z$ take values in $D(\mathcal{T},\mathbb{R})$, testing $X\operatorname*{\perp\mkern-11.0mu\perp}Y\operatorname*{|}Z$ is not straightforward. Recently, first results for such functional CI testing are proposed in Lundborg et al., (2022); Laumann et al., (2023) and Manten et al., (2024).
Example 7.
Let $\mathcal{M}_{\mathcal{D}}$ be the DSCM from Example 4. The output of FCI on the $\sigma$-separation independence model of $G((\mathcal{M}_{\mathcal{D}})_{\operatorname*{ev}^{+}(0)})$ is given in Figure 4.
[…]
Figure 4: The output of FCI on $\textrm{IM}_{\sigma}(G((\mathcal{M}_{\mathcal{D}})_{\operatorname*{ev}^{+}(0)}))$.
We conjecture that ‘Tiered FCI’ (Andrews et al., , 2020) is sound for cyclic DSCMs, and applying it on evaluated DSCM $\mathcal{M}_{\operatorname*{ev}(t)}$ and subsequently mapping its output back to the PAG for $\mathcal{M}$ can be used to orient more edges in the output of regular FCI.
7 Conclusion
In this work, we have refined the concept of a Dynamic Structural Causal Model. We have shown that a large class of systems of differential equations can be mapped to a DSCM, giving it a Markov property based on $\sigma$-separation. We have formalised the concepts of time-splitting and subsampling, which both subsequently result in a DSCM, and thus remain to have clear causal semantics. We have shown how a time-split DSCM can shed some light on the concept of local conditional independence, and that the graph of a collapsed DSCM can be interpreted as a local independence graph if all integrators are exogenous and independent. Additionally, we have shown how existing results for causal effect identification and constraint-based causal discovery for simple SCMs can directly be applied to DSCMs.
Appendix A Adaptedness of solutions of simple DSCMs
Naturally, for a solution $(X_{V},X_{W}):\Omega\times\mathcal{T}\to\mathbb{R}^{|V\cup W|}$ of a DSCM $\mathcal{M}$ and $\mathbb{F}_{X_{W}}$ the filtration on $\Omega$ consisting of sigma algebras $\mathcal{F}_{t}:=\sigma(\{X_{W}(s):s\leq t\})$, one would require $X_{V_{p}}$ to be $\mathbb{F}_{X_{W}}$-adapted. However, for intervened DSCMs $\mathcal{M}_{\operatorname*{\mathrm{do}}(X_{T}=x_{T})}$ we also want the solution to not depend on future values of $x_{T}$; a property that cannot be expressed in terms of a filtration on $\Omega$, since the intervened variable $X_{T}$ does not depend on $\Omega$. To mitigate this problem, we require adaptedness of the solution functions with respect to the canonical filtration on $\mathcal{X}_{W}$, which we will define shortly.
The classical approach for modelling stochastic processes is to consider a probability space $(\Omega,\mathcal{F},\mathbb{P})$, and models for example a Brownian motion $X:\Omega\times\mathcal{T}\to\mathbb{R}$ and a Poisson process $Y:\Omega\times\mathcal{T}\to\mathbb{R}$, for which the induced filtrations $\mathbb{F}_{X}$ and $\mathbb{F}_{Y}$ on $\Omega$ are different. We can equivalently model the exogenous processes directly on $D(\mathcal{T},\mathbb{R})$ via the laws $\mathbb{P}(X)$ and $\mathbb{P}(Y)$ (so without some underlying measurable space $\Omega$): the corresponding stochastic processes $\tilde{X}$ and $\tilde{Y}$ on $D(\mathcal{T},\mathbb{R})$ are modelled as the coordinate mapping process $Z:D(\mathcal{T},\mathbb{R})\times\mathcal{T}\to\mathbb{R},(\omega,t)\mapsto\omega(t)$, and equipping the measurable space $D(\mathcal{T},\mathbb{R})$ either with $\mathbb{P}(X)$ or $\mathbb{P}(Y)$ determines whether $Z$ is a Brownian motion or Poisson process. The canonical filtration $\mathbb{F}$ on $D(\mathcal{T},\mathbb{R})$ is the filtration generated by $Z$.
Definition 16 (Canonical filtration).
The canonical filtration on $D(\mathcal{T},\mathbb{R})$ is defined as $\mathbb{F}=\{\mathcal{F}_{t}:t\in\mathcal{T}\}$ with
$\mathcal{F}_{t}=\sigma\left(\left\{z\in D(\mathcal{T},\mathbb{R}):z(s)\in B,s\leq t,B\in\mathcal{B}(\mathbb{R})\right\}\right).$
An adapted, measurable function $Y=g(Z)$ of such a coordinate process $Z$ then generates a subfiltration $\mathcal{F}_{t}^{Y}\subseteq\mathcal{F}_{t}$. In Proposition 5, we show that solutions of simple DSCMs are adapted to the canonical filtration on the outcome spaces of the exogenous (and intervened, if applicable) variables.
Lemma 1.
For given process $X$ and $Y:=g(X)$ with $g$ measurable, we have $Y\in\mathcal{F}_{t}^{X}$ if $g$ is adapted, and $Y\in\mathcal{F}_{t-}^{X}$ if $g$ is predictable, where $\mathcal{F}_{t}^{X}$ denotes the filtration generated by $X$.
Proof.
We have
$\mathcal{F}_{t}^{X}$ $=\sigma(\{X(s):s\leq t\})$
=$=\{\pi_{s}(X)\in B\mid s\leq t,B\in\mathcal{B}(\mathbb{R})\}$
and
$\mathcal{F}_{t-}^{X}=\{\pi_{s}(X)\in B\mid s<t,B\in\mathcal{B}(\mathbb{R})\}$
similarly. Let $Z=g(X)$ with $g$ adapted. For $t\in\mathcal{T}$ we have
$\{Z(t)$ $\in B\mid B\in\mathcal{B}(\mathbb{R})\}$
=$=\{\pi_{t}\circ g\circ X\in B\mid B\in\mathcal{B}(\mathbb{R})\}$
=$=\{X\in g^{-1}(\pi_{t}^{-1}(B))\mid B\in\mathcal{B}(\mathbb{R})\}$
=$=\{\pi_{s}(X)\in\pi_{s}(g^{-1}(\pi_{t}^{-1}(B)))\mid s\leq t,B\in\mathcal{B}(\mathbb{R})\}.$
where the last equality holds since $g$ is adapted. This gives
$\{Z(t)$ $\in B\mid B\in\mathcal{B}(\mathbb{R})\}$
$\subseteq\{\pi_{s}(X)\in B\mid s\leq t,B\in\mathcal{B}(\mathbb{R})\}=\mathcal{F}_{t}^{X}.$
Similarly, if $Z=g(X)$ with $g$ predictable, we have
$\{Z(t)$ $\in B\mid B\in\mathcal{B}(\mathbb{R})\}$
=$=\{\pi_{s}(X)\in\pi_{s}(g^{-1}(\pi_{t}^{-1}(B)))\mid s<t,B\in\mathcal{B}(\mathbb{R})\}$
$\subseteq\{\pi_{s}(X)\in B\mid s<t,B\in\mathcal{B}(\mathbb{R})\}=\mathcal{F}_{t-}^{X}.$
Proposition 5 (Adaptedness of solutions).
Given an intervened simple DSCM $\mathcal{M}_{\operatorname*{\mathrm{do}}(X_{T}=x_{t})}$ (write $O:=V\setminus T$), the solution $X_{O}$ of the is adapted to the canonical filtration on $\mathcal{X}_{T\cup W}$.
Proof.
Since $\mathcal{M}_{\operatorname*{\mathrm{do}}(X_{T}=x_{t})}$ is simple there is a solution function $g_{O}$ such that $X_{O}=g_{O}(x_{T},X_{W})$. By Proposition 2 we have that $g_{O}$ is adapted, so by Lemma 1 we have that $X_{O}$ is adapted to $\mathcal{F}_{t}^{X_{T},X_{W}}$. Since we consider $(X_{T},X_{W})$ to be the coordinate process on $\mathcal{X}_{T}\times\mathcal{X}_{W}$, the filtration $\mathcal{F}_{t}^{X_{T},X_{W}}$ is the canonical filtration on $\mathcal{X}_{T}\times\mathcal{X}_{W}$. ∎
Appendix B DSCM induced by a system of SDEs
To prove that $\mathcal{M}_{\mathcal{D}}$ is simple, we use the following notion:
Definition 17 (Perfect interventions on SDEs).
Given a system of SDEs $\mathcal{D}=\left\langle V,W,(X_{V}^{0},X_{W}),f\right\rangle$, intervention target $T\subseteq V$ and intervention value $x_{T}\in D(\mathcal{T},\mathbb{R}^{|T|})$, the intervened system of SDEs is defined as $\mathcal{D}_{\operatorname*{\mathrm{do}}(X_{T}=x_{T})}=\left\langle V,W,(X_{V}^{0},X_{W}),f^{\circ}\right\rangle$ with
$f^{\circ}_{v}:\begin{cases}X_{v}(t)=X_{v}^{0}\\ \quad+\int_{0}^{t}g_{v}(s-,X_{\alpha(v)})\mathrm{d}h_{v}(s,X_{\beta(v)})&\text{if $v\in V\setminus T$}\\ X_{v}(t)=x_{v}(t)&\text{if $v\in T$},\end{cases}$
where $x_{T}$ is to be interpreted as an element of $\mathbb{S}_{F}$, so as function $x_{T}:\Omega\times\mathcal{T}\to\mathbb{R}^{|T|}$ that is constant in $\Omega$.
A definition of a soft-intervention on a system of SDEs can be found in Hansen and Sokol, (2014).
Proposition 6.
Given a system of SDEs $\mathcal{D}=\left\langle V,W,(X_{V}^{0},X_{W}),f\right\rangle$, intervention target $T\subseteq V$ and intervention value $x_{T}\in\mathcal{X}_{T}$, we have
$\mathcal{M}_{(\mathcal{D}_{\operatorname*{\mathrm{do}}(X_{T}=x_{T})})}=(\mathcal{M}_{\mathcal{D}})_{\operatorname*{\mathrm{do}}(X_{T}=x_{T})}.$
Proof.
We have $\mathcal{M}_{\mathcal{D}}=\left\langle V,W,\mathcal{X}_{V},\mathcal{X}_{W},f^{*},\mathbb{P}(X_{W})\right\rangle$ and $(\mathcal{M}_{\mathcal{D}})_{\operatorname*{\mathrm{do}}(X_{T}=x_{T})}=\left\langle V,W,\mathcal{X}_{V},\mathcal{X}_{W},(f^{*})^{\circ},\mathbb{P}(X_{W})\right\rangle$, and on the other hand $\mathcal{D}_{\operatorname*{\mathrm{do}}(X_{T}=x_{T})}=\left\langle V,W,(X_{V}^{0},X_{W}),f^{\circ}\right\rangle$ and $\mathcal{M}_{(\mathcal{D}_{\operatorname*{\mathrm{do}}(X_{T}=x_{T})})}=\left\langle V,W,\mathcal{X}_{V},\mathcal{X}_{W},(f^{\circ})^{*},\mathbb{P}(X_{W})\right\rangle$. Since the trivial SDE $X_{T}(t)=x_{T}(t)$ has Itô map $I_{T}(x_{T})=x_{T}$, it is clear that taking the Itô map and intervention commute, i.e. $(f^{*})^{\circ}=(f^{\circ})^{*}$, so we have the required result. ∎
Proposition 7.
Given a uniquely solvable system of SDEs $\mathcal{D}$, the induced DSCM $\mathcal{M}_{\mathcal{D}}$ is simple.
Proof.
For every $T\subseteq V$ and $x_{T}\in\mathcal{X}_{T}$, let $\mathcal{D}^{\prime}:=\left\langle V\setminus T,W\cup T,(X_{V\setminus T}^{0},X_{W},x_{T}),f_{V\setminus T}\right\rangle$, where $x_{T}$ is to be interpreted as an element of $\mathbb{S}_{F}$ that is constant in $\Omega$. Then $\mathcal{D}^{\prime}$ is uniquely solvable, hence it has an Itô map $I_{V\setminus T}(x_{V}^{0},x_{T},x_{W})$. Considering the intervened SDE $\mathcal{D}_{\operatorname*{\mathrm{do}}(X_{T}=x_{T})}$ defined according to Definition 17, we have that $I_{V\setminus T}(x_{V}^{0},x_{T},x_{W})$ is the solution function of $\mathcal{M}_{(\mathcal{D}_{\operatorname*{\mathrm{do}}(X_{T}=x_{T})})}$. By Proposition 6 we have that $I_{V\setminus T}$ is the solution function of $(\mathcal{M}_{\mathcal{D}})_{\operatorname*{\mathrm{do}}(X_{T}=x_{T})}$ as well, so we have unique solvability of $\mathcal{M}_{\mathcal{D}}$. ∎
Appendix C Proofs
Proposition 1 (Unique solvability of a system of SDEs).
Let $\mathcal{D}=\left\langle V,W,(X_{V}^{0},X_{W}),f\right\rangle$ be a system of SDEs such that for every SDE $f_{v}$, the parameters $g_{v}$ and $h_{v}$ satisfy Assumptions 2 and 1 respectively. If for every $v\in V$ we have $\beta(v)\cap\operatorname*{\mathrm{Sc}}(v)=\emptyset$, there exists an Itô map $I_{V}$ such that $X_{V}:=I_{V}(X_{V}^{0},X_{W})\in\mathbb{S}_{\mathbb{F}}(\mathcal{T},\mathbb{R}^{|V|})$ is the unique solution of $\mathcal{D}$.
Proof.
Let $(S_{1},...,S_{k})$ with $k\leq|V|$ be a topologically ordered partition of $V$ into strongly connected components of $G(\mathcal{D})$. For the first component $S_{1}$ write $\tilde{\operatorname*{\mathrm{Pa}}}(S_{1})=\alpha(S_{1})\cup\beta(S_{1})\setminus S_{1}$, which has $\tilde{\operatorname*{\mathrm{Pa}}}(S_{1})\subseteq W$, hence we can apply Theorem 1 to obtain $X_{S_{1}}:=I_{S_{1}}(X_{S_{1}}^{0},X_{\tilde{\operatorname*{\mathrm{Pa}}}(S_{1})})$ as the solution of the SDE $f_{S_{1}}$. For given $i\in\{1,...,k\}$ denote $S_{\leq i}:=\bigcup_{j\leq i}S_{j}$, and assume that $X_{S_{\leq i}}:=I_{S_{\leq i}}(X_{S_{\leq i}}^{0},X_{\tilde{\operatorname*{\mathrm{Pa}}}(S_{\leq i})\cap W})$ is the solution of the SDE $f_{S_{\leq i}}$. We have that $\tilde{\operatorname*{\mathrm{Pa}}}(S_{i+1})\subseteq S_{\leq i}\cup W$. We apply Theorem 1 again to obtain the Itô map $I_{S_{i+1}}(X_{S_{i+1}}^{0},X_{\tilde{\operatorname*{\mathrm{Pa}}}(S_{i+1})\cap S_{\leq i}},X_{\tilde{\operatorname*{\mathrm{Pa}}}(S_{i+1})\cap W})$. The Itô map for the SDE $f_{S_{\leq i+1}}$ is then $I_{S_{\leq i+1}}(X_{V}^{0},X_{W})$ with as its $S_{i+1}$-component $I_{S_{i+1}}(X_{S_{i+1}}^{0},I_{\tilde{\operatorname*{\mathrm{Pa}}}(S_{i+1})\cap V}(X_{V}^{0},X_{W}),X_{\tilde{\operatorname*{\mathrm{Pa}}}(S_{i+1})\cap W})$, which gives a unique solution $X_{S_{\leq i+1}}:=I_{S_{\leq i+1}}(X_{V}^{0},X_{W})$. Finally, set $I_{V}:=I_{S_{\leq k}}$. ∎
Proposition 2.
Given a simple DSCM $\mathcal{M}$, the solution function $g_{O}(x_{T},x_{W})$ of the intervened DSCM $\mathcal{M}_{\operatorname*{\mathrm{do}}(X_{T}=x_{T})}$ is adapted.
Proof.
By definition of the solution function $g_{O}$ and adaptedness of $f_{O}$ we have for all $x_{T}\in\mathcal{X}_{T}$ and $\mathbb{P}(X_{W})$-almost all $x_{W}\in\mathcal{X}_{W}$ that
$g_{O}(x_{T},x_{W})^{t}$ $=f_{O}(g_{O}(x_{T},x_{W}),x_{T},x_{W})^{t}$
$=f_{O}(g_{O}(x_{T},x_{W})^{t},x_{T}^{t},x_{W}^{t})^{t}$
$=f_{O}(g_{O}(x_{T}^{t},x_{W}^{t})^{t},x_{T}^{t},x_{W}^{t})^{t}$
$=g_{O}(x_{T}^{t},x_{W}^{t})^{t},$
so $g_{O}$ is adapted. ∎
Proposition 3.
The class of simple DSCMs is closed under marginalisation.
Proof.
By Bongers et al., (2021), Proposition 8.2 we have that $\mathcal{M}_{\operatorname*{\mathrm{marg}}(L)}$ is a simple SCM. Since $\tilde{f}(x_{O},x_{W})^{t}=f_{O}(x_{O},g_{L}(x_{O},x_{W}),x_{W})^{t}=f_{O}(x_{O}^{t},g_{L}(x_{O}^{t},x_{W}^{t})^{t},x_{W}^{t})^{t}=\tilde{f}(x_{O}^{t},x_{W}^{t})^{t}$ holds by adaptedness of $f$ and $g_{L}$, we also have that its structural equation $\tilde{f}$ is adapted, and so $\mathcal{M}_{\operatorname*{\mathrm{marg}}(L)}$ is indeed a simple DSCM. ∎
Proposition 4.
For a DSCM $\mathcal{M}$ with evaluation index $\mathcal{E}$ that partitions $\mathcal{T}$ and a sequence of increasing time-indices $\tau=(t_{1},...,t_{m})\subseteq\mathcal{T}$, we have $\mathcal{M}=\mathcal{M}_{\operatorname*{ev}^{+}(\tau)*}$. If $\tau\neq\mathcal{T}$, we have $\mathcal{M}\neq\mathcal{M}_{\operatorname*{ev}(\tau)*}$.
Proof.
The first claim follows directly from Definition 13. For the second claim, let $V\times\mathcal{E}$ and $V\times\mathcal{E}^{\prime}$ denote the endogenous indices of $\mathcal{M}$ and $\mathcal{M}_{\operatorname*{ev}(\tau)*}$ respectively. If $\tau\neq\mathcal{T}$ we have $\mathcal{E}^{\prime}\neq\mathcal{E}$, which proves the result. ∎
C.1 Local independence
As a link between the notion of local independence and conditional independence of the sample paths of stochastic processes, we relate a conditional distribution $\mathbb{E}[X_{B}^{t}\mid\mathcal{F}_{t}^{A}]:\Omega\to\mathbb{R}$ to the conditional distribution $\mathbb{E}\left[X_{B}^{t}\mid X_{A}^{[0,t]}\right]:\mathbb{R}\to\mathbb{R}$.
Lemma 2.
Given a DSCM $\mathcal{M}$ with $A,B\subseteq V$, in $\mathcal{M}_{\operatorname*{ev}(t)}$ we have $\mathbb{E}\left[X_{B}^{t}\mid\mathcal{F}^{A}_{t}\right]=\mathbb{E}\left[X_{B}^{t}\mid X_{A}^{[0,t]}=X_{A}^{[0,t]}\right]$ a.s. and $\mathbb{E}\left[X_{B}^{t}\mid\mathcal{F}^{A}_{t-}\right]=\mathbb{E}\left[X_{B}^{t}\mid X_{A}^{[0,t)}=X_{A}^{[0,t)}\right]$ a.s.
Proof.
First, considering the random variable $X_{A}^{[0,t]}:D(\mathcal{T},\mathbb{R}^{|W|})\to D([0,t],\mathbb{R}^{|A|})$, we have
$\sigma(X_{A}^{[0,t]}(s):s\leq t)$
$=\sigma(\{\{x_{W}:g_{A}^{[0,t]}(x_{W})(s)\in F\}\mid s\leq t,F\in\mathcal{B}(\mathbb{R})\})$
$=\sigma(\{\{x_{W}:g_{A}^{[0,t]}(x_{W})\in F^{\prime}\}\mid F^{\prime}\in\mathcal{B}(D([0,t],\mathbb{R}^{|A|}))\})$
$=\sigma(X_{A}^{[0,t]}),$
since $\mathcal{B}(D([0,t],\mathbb{R}))=\sigma(\omega_{s}:s\in[0,t])=\sigma(\{\{x_{A}:x_{A}(s)\in F\}:s\in[0,t],F\in\mathcal{B}(\mathbb{R})\})$. We then have
$\mathbb{E}\left[X_{B}^{t}\mid\mathcal{F}^{A}_{t}\right]$ $=\mathbb{E}\left[X_{B}^{t}\mid\sigma(X_{A}^{[0,t]}(s):s\leq t)\right]$
$=\mathbb{E}\left[X_{B}^{t}\mid\sigma(X_{A}^{[0,t]})\right]$
$=\mathbb{E}\left[X_{B}^{t}\mid X_{A}^{[0,t]}=X_{A}^{[0,t]}\right],$
where the last equality holds by the factorisation lemma (Kallenberg, (2021), Lemma 1.14). We also have
$\mathcal{F}_{t-}^{A}=\sigma\left(\mathcal{F}_{s}^{A}:s<t\right)=\sigma(X_{A}^{[0,t]}(s):s<t)=\sigma(X_{A}^{[0,t)}),$
and so by the factorisation lemma $\mathbb{E}\left[X_{B}^{t}\mid\mathcal{F}^{A}_{t-}\right]=\mathbb{E}\left[X_{B}^{t}\mid X_{A}^{[0,t)}=X_{A}^{[0,t)}\right]$. ∎
Theorem 3.
Considering DSCM $\mathcal{M}$ and for every $t\in\mathcal{T}$ the time-split DSCM $\mathcal{M}_{\operatorname*{ev}^{+}(t)}$, we have for all $A,B,C\subseteq V$ that
$X_{B}^{t}\operatorname*{\perp\mkern-11.0mu\perp}X_{A}^{[0,t)}|X_{B}^{[0,t)},X_{C}^{[0,t)}\text{ for all $t\in\mathcal{T}$}$
$\implies$ $X_{A}\not\to X_{B}\mid X_{C}$
$\implies$ $\mathbb{E}[X_{B}^{t}\mid X_{A}^{[0,t)},X_{B}^{[0,t)},X_{C}^{[0,t)}]=\mathbb{E}[X_{B}^{t}\mid X_{B}^{[0,t)},X_{C}^{[0,t)}]$
for all $t\in\mathcal{T}$.
Proof.
First, by Lemma 2 we have that $\mathbb{E}[X_{B}(t)\mid X_{A}^{[0,t)},X_{B}^{[0,t)},X_{C}^{[0,t)}]=\mathbb{E}[X_{B}(t)\mid X_{B}^{[0,t)},X_{C}^{[0,t)}]$ holds a.s. iff $\mathbb{E}\left[X_{B}(t)\mid\mathcal{F}^{A,B,C}_{t-}\right]=\mathbb{E}\left[X_{B}(t)\mid\mathcal{F}^{B,C}_{t-}\right]$ a.s. By the innovation theorem (see e.g. Aalen, (1978)), we can pick a version $\lambda^{B,C}(t)$ of $\mathbb{E}[\lambda^{A,B,C}(t)\mid\mathcal{F}_{t-}^{B,C}]$ that is a $\mathcal{F}_{t}^{B,C}$-intensity of $X_{B}$. By assumption, we can write
$X_{B}(t)$ $=\int_{0}^{t}\lambda^{A,B,C}(s)\mathrm{d}s+M^{A,B,C}(t)$
$=\int_{0}^{t}\lambda^{B,C}(s)\mathrm{d}s+M^{B,C}(t)$
for $M^{A,B,C}$ ($M^{B,C}$) a $\mathcal{F}_{t}^{A,B,C}$ (resp. $\mathcal{F}_{t}^{B,C}$) martingale.
By Aalen, (1978), we have $\lambda_{t}^{A,B,C}=\lim_{h\downarrow 0}\frac{1}{h}\mathbb{E}[X_{B}(t+h)-X_{B}(t)\mid\mathcal{F}_{t-}^{A,B,C}]$. Since $\lambda^{A,B,C}$ and $\lambda^{B,C}$ are bounded, by dominated convergence we have $\lambda_{t}^{A,B,C}=\mathbb{E}[\lim_{h\downarrow 0}\frac{1}{h}(X_{B}(t+h)-X_{B}(t))\mid\mathcal{F}_{t-}^{A,B,C}]$. Since $\lim_{h\downarrow 0}\frac{1}{h}(X_{B}(t+h)-X_{B}(t))\in\mathcal{F}_{t}^{B}$ and the independence assertion can be equivalently stated as $\mathcal{F}_{t}^{B}\operatorname*{\perp\mkern-11.0mu\perp}\mathcal{F}_{t-}^{A}|\mathcal{F}_{t-}^{B,C}$, we get
$\lambda_{t}^{A,B,C}$ $=\mathbb{E}\left[\lim_{h\downarrow 0}\frac{1}{h}(X_{B}(t+h)-X_{B}(t))\mid\mathcal{F}_{t-}^{A,B,C}\right]$
$=\mathbb{E}\left[\lim_{h\downarrow 0}\frac{1}{h}(X_{B}(t+h)-X_{B}(t))\mid\mathcal{F}_{t-}^{B,C}\right]$
$=\lambda^{B,C}_{t},$
proving local independence.
Under local conditional independence we have $\lambda^{A,B,C}=\lambda^{B,C}$ a.s., and by the above equation also $M^{A,B,C}=M^{B,C}$ a.s., so by Jacod and Shiryaev, (2003), Lemma I.2.27 we have that
$\mathbb{E}\left[M^{A,B,C}(t)|\mathcal{F}_{t-}^{A,B,C}\right]=M^{A,B,C}(t-)\\ =M^{B,C}(t-)=\mathbb{E}\left[M^{B,C}(t)|\mathcal{F}_{t-}^{B,C}\right]$
and so
$\mathbb{E}\left[X_{B}(t)\mid\mathcal{F}^{A,B,C}_{t-}\right]$
$=\mathbb{E}\left[\int_{0}^{t}\lambda^{A,B,C}(s)\mathrm{d}s+M^{A,B,C}(t)\mid\mathcal{F}^{A,B,C}_{t-}\right]$
$=\mathbb{E}\left[\int_{0}^{t}\lambda^{B,C}(s)\mathrm{d}s+M^{B,C}(t)\mid\mathcal{F}^{B,C}_{t-}\right]$
$=\mathbb{E}\left[X_{B}(t)\mid\mathcal{F}^{B,C}_{t-}\right],$
so $\mathbb{E}\left[X_{B}(t)\mid X_{A}^{[0,t)},X_{B}^{[0,t)},X_{C}^{[0,t)}\right]=\mathbb{E}\left[X_{B}(t)\mid X_{B}^{[0,t)},X_{C}^{[0,t)}\right]$. ∎
Theorem 4.
For a collapsed DSCM $\mathcal{M}$ that satisfies Assumption 3 we have:
  1. 1.
    $G(\mathcal{M})$ is a local independence graph,
  2. 2.
    a $\sigma$-separation local independence global Markov property:
    $A\operatorname*{\perp}_{G(\mathcal{M})}^{\sigma}B|C\implies X_{A}\not\to X_{B}\mid X_{C}$
    for all $A,B,C\subseteq V$.
Proof.
Throughout this proof, write $G_{t}:=G(\mathcal{M}_{\operatorname*{ev}^{+}(t)})$.
1) Let $A,B\in V$. By Assumption 3 we have that $\operatorname*{\mathrm{Pa}}_{G(\mathcal{M})}(B)=\alpha(B)$. If $A\to B$ not in $G(\mathcal{M})$, we have $A\notin\alpha(B)$, so $X_{A}^{[0,t)}\notin\operatorname*{\mathrm{Pa}}_{G_{t}}(X_{B}^{t})=X_{\alpha(B)}^{[0,t)}$ for all $t$. If we let $N:=(V\setminus A)\setminus\alpha(B)$, then since $X_{A}^{[0,t)}\in\operatorname*{\mathrm{Nd}}(X_{B}^{t})$ and $X_{N}^{[0,t)}\subset\operatorname*{\mathrm{Nd}}(X_{B}^{t})$ we have by the local Markov property [^14] For SCM $\mathcal{M}$ with $B\in V$, we have $X_{B}\operatorname*{\perp\mkern-11.0mu\perp}X_{\operatorname*{\mathrm{Nd}}(B)}|X_{\operatorname*{\mathrm{Pa}}(B)}$, where $\operatorname*{\mathrm{Nd}}(B)=V\setminus\operatorname*{\mathrm{Desc}}(B)$ is the set of non-descendants of $B$. that $X_{B}^{t}\operatorname*{\perp\mkern-11.0mu\perp}X_{A}^{[0,t)}|X_{\alpha(B)}^{[0,t)}$ and $X_{B}^{t}\operatorname*{\perp\mkern-11.0mu\perp}X_{N}^{[0,t)}|X_{\alpha(B)}^{[0,t)}$. By the weak contraction property of conditional independence (and the fact that $V\setminus A=\alpha(B)\cup N$) this gives $X_{B}^{t}\operatorname*{\perp\mkern-11.0mu\perp}X_{A}^{[0,t)}|X_{V\setminus A}^{[0,t)}$ so by Theorem 3 we have $X_{A}\not\to_{\mathbb{P}}X_{B}|X_{V\setminus A}$.
2) Let $G^{\prime}$ be the $[0,t)$-slice of $G_{t}$, so $G^{\prime}:=G((\mathcal{M}_{\operatorname*{ev}^{+}(t)})_{\operatorname*{\mathrm{marg}}(V\times(\mathcal{E}\setminus\{[0,t)\}))})$, for which we have $G^{\prime}\subseteq G(\mathcal{M})$, hence $X_{B}^{[0,t)}\operatorname*{\perp}_{G^{\prime}}^{\sigma}X_{A}^{[0,t)}|X_{C}^{[0,t)}$. Since all vertices in $G^{\prime}$ have no additional ancestors in $G_{t}$, we have $X_{B}^{[0,t)}\operatorname*{\perp}_{G_{t}}^{\sigma}X_{A}^{[0,t)}|X_{C}^{[0,t)}$. We have that every walk $\pi$ from $X_{A}^{[0,t)}$ to $X_{B}^{t}$ that is $\sigma$-open given $X_{C}^{[0,t)}$ must end in $v\to X_{B}^{t}$ for some $v\in\{X_{B}^{[0,t)}\}\cup\operatorname*{\mathrm{Pa}}(X_{B}^{[0,t)})$, hence a copy $\pi^{\prime}$ of that walk that ends in $v\to X_{B}^{[0,t)}$ must be $\sigma$-open given $X_{C}^{[0,t)}$ as well, contradicting $X_{B}^{[0,t)}\operatorname*{\perp}_{G_{t}}^{\sigma}X_{A}^{[0,t)}|X_{C}^{[0,t)}$, so we must have $X_{B}^{t}\operatorname*{\perp}_{G_{t}}^{\sigma}X_{A}^{[0,t)}|X_{C}^{[0,t)}$ for all $t$. Since this implies $X_{B}^{t}\operatorname*{\perp}_{G_{t}}^{\sigma}X_{A}^{[0,t)}|X_{B}^{[0,t)},X_{C}^{[0,t)}$, Theorem 3 gives the result. ∎